OpenStudy (anonymous):

The perimeter of a triangle is 48 inches and the sides are in the ratio of 3:4:5. Find the length of the longest side. A.4 B.9 C.16 D.20

OpenStudy (anonymous):

3x+4x+5x=48 12x=48 x=48/12 x=4 substitute the x=4 into these equations to get the longest side: 3x=3(4)=12 4x=3(4)=16 5x=5(4)=20 therefore the longest side is 20.

OpenStudy (anonymous):

Oh thank you that makes it a lot easier I think I get it now

OpenStudy (anonymous):

your welcome, hope that helps alot

OpenStudy (raden):

alternative : the longest side = 5/(3+4+5) * perimeter = 5/12 * 48 = ...